### v1

# On the closure of a plane ray that limits onto itself

Abstract.
We show that the closure of any self-entwined ray in the plane must contain an uncountable collection of pairwise-disjoint subcontinua. This is false in dimension three.
Key words and phrases:
self-entwined, non-Suslinian, ray, indecomposable, plane topology
2020 Mathematics Subject Classification:
54F15, 37B45
1. Introduction
In this paper we investigate properties of planar continua which contain self-entwined rays.
1.1. Terminology
A continuum is a compact connected metric space. A ray is a continuous one-to-one image of the interval $0,\infty)$. A ray is self-entwined, or is a ray limiting onto itself, if no open set intersects it in an arc (see Section 2.2). A continuum is non-Suslinian if it contains an uncountable collection of pairwise-disjoint, non-degenerate subcontinua. A connected space $X$ is decomposable if there exist proper closed connected subsets $A$ and $B$ such that $X=A\cup B$; otherwise $X$ is indecomposable. Indecomposable continua are known to be non-Suslinian.
1.2. Motivation and main result
Self-entwined rays are commonly found in continua which arise from folding algorithms and chaotic systems. Such examples exist within the arc components of the Smale horseshoe, and Ikeda and Plykin attractors (see Figure 1). These continua are clearly non-Suslinian as they are indecomposable and even have neighborhoods homeomorphic to the Cantor fence $2^{\omega}\times[0,1]$.
The goal of this paper is to establish the general result:
Theorem 1.
If $X$ is a self-entwined ray in the plane, then $\overline{X}$ is non-Suslinian.
Although every ray is a rational space, meaning that it has a basis of open sets with countable boundaries, Theorem 1 implies that the closure of a self-entwined plane ray cannot be rational.
1.3. Context and limitations
Jones [5, 6] had shown that a plane ray cannot be locally connected unless it is locally compact. Curry [2] proved that the closure of a self-entwined plane ray is an indecomposable continuum provided that it is $1$-dimensional and non-separating or finitely separating. On the other hand, it can be shown that the Sierpiński carpet, which is locally connected and $1$-dimensional, contains a dense self-entwined ray. There even exists a self-entwined ray in the plane that has $1$-dimensional closure, and is itself a decomposable connected set [9]. This example shows that it is not always possible to “open up” complementary components of the closure and derive the non-Suslinian property from Curry’s result (cf. [4, Section 2.1]).
Theorem 1 fails in higher dimensions by an example of Tymchatyn [13, Example 3]. Tymchatyn’s continuum was already known to be non-planar because it is hereditarily locally connected but not finitely-Suslinian [7].
Figure 1. An Ikeda attractor. From [14]: “The Ikeda map is a discrete-time dynamical system that produces a strange attractor. It was introduced in 1979 by the physicist Kensuke Ikeda as a model for the behavior of light within a nonlinear optical resonator. The map demonstrates how a simple set of rules can lead to complex, chaotic behavior through a process of repeated rotation, scaling, and translation—a ‘stretch and fold’ operation common in chaotic systems.”
2. Preliminaries
2.1. Arcs of rays
An arc is a topological copy of the interval $[0,1]$. Note that if $X$ is any ray with an associated mapping $f$, and $a,b]\subset[0,\infty)$ is a closed and bounded interval, then $f\restriction[a,b]$ is a homeomorphism and $f[a,b]$ is an arc in $X$. It is a consequence of Sierpiński’s theorem [12, Theorem 5.16] that if $X$ is not compact then every continuum in $X$ is of this form [1, Lemma 2.2].
2.2. Definition of self-entwined
Here is an equivalent formulation of the self-entwined property in terms of the ray function.
Proposition 2.
If $X$ is a ray and $f$ is a continuous one-to-one mapping of $0,\infty)$ onto $X$, then the following are equivalent.
  • Each arc of $X$ has empty interior in $X$;
  • For every $x\in X$ there exists a sequence $r_{n}\in[0,\infty)$ such that $r_{n}\to\infty$ and $f(r_{n})\to x$.
Other equivalent formulations include that the space $X$ has no compact neighborhoods, or is first category in the sense of Baire.
2.3. Connectedness im-kleinen
A space $X$ is said to be connected im-kleinen at a point $x\in X$ if every neighborhood of $x$ contains a connected neighborhood of $x$.
The following is due to Fitzpatrick and Lelek [3].
Proposition 3.
Let $X$ be a continuum. If $X$ is Suslinian, then $X$ is connected im-kleinen at each point of a dense subset.
Idea of proof.
If $U$ is a non-empty open subset of $X$ that has only countably many non-degenerate components, then some component of $U$ must have interior in $X$ by the Baire category theorem. So if $X$ is Suslinian, it is possible to recursively define a decreasing sequence of continua with interiors in $X$, which intersect down to a single point of $X$ that is contained in each interior. ∎
2.4. Discs and theta curves
A $\theta$-curve is a space made of three arcs $\alpha_{1},\alpha_{2},\alpha_{3}$ which have the same endpoints $a$ and $b$ and are otherwise disjoint. It is a standard fact of plane topology that one of the three arcs comprising a $\theta$-curve must be separated from $\infty$ by the union of the other two.
Proposition 4.
Let $\theta=\alpha_{1}\cup\alpha_{2}\cup\alpha_{3}$ be a theta curve of arcs joining $a$ and $b$, such that $\alpha_{1}\cup\alpha_{3}$ separates $\alpha_{2}$ from $\infty$. Let $D$ be a closed disc missing $a$ and $b$. If $p\in\alpha_{1}\cap D$ and $q\in\alpha_{3}\cap D$, then there exists an arc $\beta\subset\alpha_{2}\cap D$ which separates $p$ from $q$ in $D$, and is such that $\beta\cap\partial D$ consists only of the two endpoints of $\beta$.
Sketch of proof.
Let $\alpha_{4}$ be an arc from $a$ to $b$ in the unbounded component of $\mathbb{R}^{2}\setminus(\alpha_{1}\cup\alpha_{3})$ that misses $D$. Then the simple closed curve $\alpha_{2}\cup\alpha_{4}$ separates $p$ and $q$. It follows that $\alpha_{2}\cap D$ separates $p$ from $q$ in $D$. Let $C$ be the component of $p$ in $D\setminus\alpha_{2}$. Then $\partial C$ is a simple closed curve and $\partial C\cap\alpha_{2}$ is an arc meeting the interior of $D$. Choose $x\in\alpha_{2}\cap\partial C\cap D^{\mathrm{o}}$ and let $\beta$ be the closure of the component of $x$ in $\partial C\cap D^{\mathrm{o}}$. ∎
2.5. Notation
For every $A\subset\mathbb{R}^{2}$ we let $\overline{A}$, $\partial A$, and $A^{\mathrm{o}}$ denote the closure, boundary, and interior of $A$, respectively, relative to $\mathbb{R}^{2}$.
We denote by $2^{n}$ the set of functions from $\{0,1,\ldots,n-1\}$ to $\{0,1\}$. Thus each $\sigma\in 2^{n}$ can be viewed as a binary sequence of length $n$. That is, $\sigma=\langle\sigma(0),\ldots,\sigma(n-1)\rangle$. We let
$2^{<\omega}=\bigcup_{n=0}^{\infty}2^{n}$
denote the set of all finite binary sequences, and finally $2^{\omega}$ is the set of all infinite binary sequences index by $\omega=\{0,1,2,\ldots\}$.
If $\sigma\in 2^{<\omega}$, then $\sigma^{\frown}0$ and $\sigma^{\frown}1$ represent the one-element extensions of $\sigma$ obtained by appending 0 and $1$, respectively. If $\tau\in 2^{\leq\omega}$ and $n<\text{dom}(\tau)$, then $\tau\restriction n$ is the sequence made of the first $n$ terms of $\tau$.
Figure 2. Illustration for Proposition 4.
3. Rays that limit to arcs on both sides
Let $X$ be a ray in $\mathbb{R}^{2}$, and let $f$ be a continuous one-to-one mapping of $0,\infty)$ onto $X$. For each integer $n\geq 1$ let $S_{n}$ be a simple closed curve containing the arc $f[0,n]$. Let $U_{n}$ and $V_{n}$ be the components of $\mathbb{R}^{2}\setminus S_{n}$.
Given $t>0$ and an integer $n>t$, we say that $X$ limits to $f(t)$ from both sides of $f[0,n]$ if $f(t)\in\overline{U_{n}\cap X}\cap\overline{V_{n}\cap X}$. A point $x=f(t)\in X$ is a two-sided limit point of $X$ if there exists an integer $n>t$ such that $X$ limits to $f(t)$ from both sides of $f[0,n]$.
Let $A,B,C\subset\mathbb{R}^{2}$. An arc $\alpha\subset\mathbb{R}^{2}$ is $ABC$-minimal if $\alpha\cap A$ and $\alpha\cap C$ are the endpoints of $\alpha$, and $\alpha\cap B\neq\varnothing$. It is easy to prove that if $A$, $B$, and $C$ are closed and pairwise-disjoint, and $\alpha$ is any arc meeting all three sets, then $\alpha$ contains an arc which is $ABC$, BACBAC, or $ACB$-minimal.
Lemma 5.
If $X$ is self-entwined and $\overline{X}$ is connected im-kleinen at a dense set of points, then the set of two-sided limit points of $X$ is dense in $X$.
Proof.
Suppose that $X$ is self-entwined and $\overline{X}$ is connected im-kleinen at a dense set of points. Let $U$ be any non-empty open subset of $\mathbb{R}^{2}$ meeting $X$. Let $D_{1},D_{2},D_{3}$ be pairwise-disjoint closed discs in $U$ whose interiors intersect $\overline{X}$. For each $i\in\{1,2,3\}$ let $K_{i}\subset X$ be a continuum which has interior in $X$ and is contained in the interior of $D_{i}$. There are 7 pairwise-disjoint arcs in $X$ which meet every $K_{i}$. Each one contains an arc that is $K_{1}K_{2}K_{3}$, or $K_{2}K_{1}K_{3}$, or $K_{2}K_{3}K_{1}$-minimal. At least 3 of the 7 minimal arcs, say $\alpha_{1},\alpha_{2},\alpha_{3}$, are of the same type. By re-indexing the $D^{\prime}s$ and $K^{\prime}s$ we may assume that the three arcs are $K_{1}K_{2}K_{3}$-minimal. Upon identifying $K_{1}$ and $K_{3}$ with points $a$ and $b$, we have that $\alpha_{1}\cup\alpha_{2}\cup\alpha_{3}$ is a $\theta$-curve. Let us assume that $\alpha_{2}$ is separated from $\infty$ by $\alpha_{1}\cup\alpha_{3}$. Choose $p\in\alpha_{1}\cap K_{2}$ and $q\in\alpha_{3}\cap K_{2}$. By Proposition 4, there exists an arc $\beta\subset\alpha_{2}\cap D_{2}$ which separates $p$ from $q$ in $D_{2}$, and is such that $\beta\cap\partial D_{2}$ consists of only the two endpoints of $\beta$. We claim that $\beta$ contains a two-sided limit point of $X$.
Let $n$ be an integer such that $\beta\subset f[0,n]$. Since $S_{n}\cap X$ is nowhere dense in $X$, $\beta$ is contained in the union of $\overline{U_{n}\cap X}$ and $\overline{V_{n}\cap X}$. If each of these sets intersects $\beta$ then by connectedness of $\beta$ there exists $x\in\overline{U_{n}\cap X}\cup\overline{V_{n}\cap X}\cap\beta$. Then $x$ is a two-sided limit point of $X$. Otherwise, one of the two sets misses $\beta$ which we will show leads to a contradiction. Without loss of generality, assume that $\overline{V_{n}\cap X}\cap\beta=\varnothing$. Let $G$ be the component of $\beta$ in $D_{2}\cap\overline{V_{n}}$. Then $G$ is a closed topological disc, $\beta\subset\partial G$, and $G\setminus\overline{V_{n}\cap X}$ is a neighborhood of $\beta$ in the space $G$. Thus there exists an arc $\gamma\subset G\setminus\overline{V_{n}\cap X}$ that has the same endpoints as $\beta$ and is otherwise contained in $G^{\mathrm{o}}$. Choosing $\gamma$ close to $\beta$ will ensure that it separates $D_{2}$ between $p$ and $q$. Moreover, $\gamma\cap\overline{X}\subset\partial D_{2}$ by the fact $(G^{\mathrm{o}}\setminus\overline{V_{n}\cap X})\cap\overline{X}=\varnothing$. Therefore $\gamma$ misses $K_{2}$, which contradicts that $K_{2}$ is connected. ∎
Remark 1.
If $\overline{X}$ is nowhere connected im-kleinen, the set of two-sided limit points of $X$ may be empty. Such is the case when $X$ is the accessible composant (arc component of the endpoint) of the buckethandle continuum.
4. Proof of Theorem 1
Let $X$ be a self-entwined ray in $\mathbb{R}^{2}$. Let $f\mathrel{\mathop{\ordinarycolon}}[0,\infty)\to X$ be a continuous one-to-one surjection. We want to show that $\overline{X}$ is non-Suslinian. This follows immediately from Proposition 3 if $\overline{X}$ is not connected im-kleinen at a dense set of points. For the remainder of the proof we may therefore assume that: $\overline{X}$ is connected im-kleinen at a dense set of points.
There exists a proper subcontinuum $K$ of $\overline{X}$ with interior in $\overline{X}$, and a disc $D\subset\mathbb{R}^{2}$ such that $D$ lies in the unbounded component of $\mathbb{R}^{2}\setminus K$ and $D^{\mathrm{o}}\cap X\neq\varnothing$. We will show that $\overline{X}$ contains an uncountable collection of pairwise-disjoint subcontinua stretching from $K$ to $D$.
By working in the unbounded component of $\mathbb{R}^{2}\setminus K$, we may assume that if $\alpha$ is any $KDK$-minimal arc, then $\mathbb{R}^{2}\setminus(K\cup\alpha)$ has exactly two complementary components. We may also assume that the ray’s endpoint $f(0)$ belongs to $K$, because there exists $f(t)\in K$ and $f[t,\infty)$ is a self-entwined ray with the same closure as $X$. Under this assumption, for every $x\in X\setminus K$ there exists a (unique) $K\{x\}K$-minimal arc in $X$. This arc is found by tracing $x$ to $K$, backwards and forwards along the ray.
For each $x\in X\setminus K$ we may now define:
  • $\alpha(x)$ to be the unique $K\{x\}K$-minimal arc of $X$,
  • $U(x)$ to be the bounded component of $\mathbb{R}^{2}\setminus(K\cup\alpha(x))$, and
  • $V(x)$ to be the unbounded component of $\mathbb{R}^{2}\setminus(K\cup\alpha(x))$.
Remark 2.
Note that $\alpha(x)\subset\partial U(x)\subset\alpha(x)\cup K$ and likewise for $V(x)$. And $\{U(x),\alpha(x),V(x)\}$ partitions the family of all $K(X\setminus K)K$-minimal arcs of $X$, with the possible exception that $\alpha(x)$ may share an endpoint in $K$ with an arc in $U(x)$ or $V(x)$.
Observation 6.
If $x$ is any two-sided limit point of $X$ in $X\setminus K$, then $x\in\overline{U(x)\cap X}\cap\overline{V(x)\cap X}.$
Proof.
Suppose that $x=f(t)$ is a two-sided limit point in $X\setminus K$. Let $n>t$ be such that $X$ limits to $x$ from both sides of $f[0,n]$. There exists a closed domain $G$ such that $x\in G^{\mathrm{o}}$,
$G$ $\cap K=\varnothing,$
$G$ $\cap S_{n}\subset f(0,n)\text{, and}$
$G$ $\cap\alpha(x)=G\cap f[0,n]=\beta,$
where $\beta$ is an arc that contains $x$ and cuts $G$ into exactly two pieces. This is easy to see if one uses the fact that $f[0,n]$ is ambiently homeomorphic to a straight arc in which $f[0,n]\cap\alpha(x)$ is a neighborhood of $x$. Let $G_{0}$ and $G_{1}$ be the components of $G\setminus\beta$. Since $G\setminus\beta\subset\mathbb{R}^{2}\setminus(K\cup\alpha(x))$ and $U(x)$ and $V(x)$ each intersect $G$, without loss of generality we may assume $G_{0}\subset U(x)$ and $G_{1}\subset V(x)$. Likewise, since $G\setminus\beta\subset\mathbb{R}^{2}\setminus S_{n}$ and $U_{n}$ and $V_{n}$ each intersect $G$, there exists $i\in\{0,1\}$ such that $G_{i}\subset U_{n}$ and $G_{1-i}\subset V_{n}$. Then $G\cap U_{n}=G_{i}$ and $G\cap V_{n}=G_{1-i}$. Now let $W$ be any open set containing $x$. Since $X$ limits to $x$ from both sides of $f[0,n]$, $W\cap G_{i}\cap X$ and $W\cap G_{1-i}\cap X$ are each non-empty. Therefore $W\cap U(x)\cap X$ and $W\cap V(x)\cap X$ are each non-empty. This shows that $x\in\overline{U(x)\cap X}\cap\overline{V(x)\cap X}$. ∎
We will now define a Cantor tree of compact regions of $\mathbb{R}^{2}$,
$\big{\{}D_{\sigma}\mathrel{\mathop{\ordinarycolon}}\sigma\in 2^{<\omega}\big{\}},$
and points $x_{\sigma}\in X$ so that:
  1. (1)
    $\alpha(x)\subset D_{\sigma}$ for all $x\in D_{\sigma}\cap X$,
  2. (2)
    $x_{\sigma^{\frown}i}\in(D\cap D_{\sigma})^{\mathrm{o}}$ for each $i\in\{0,1\}$, and
  3. (3)
    $\overline{U(x_{\sigma^{\frown}1})}\setminus K\subset U(x_{\sigma^{\frown}0})$.
First step: To begin, let $a\in D^{\mathrm{o}}\cap X$ be a two-sided limit point of $X$ provided by Lemma 5. By Observation 6 and Lemma 5 there exist two-sided limit points $b\in D^{\mathrm{o}}\cap U(a)\cap X$, $c\in D^{\mathrm{o}}\cap U(b)\cap X$, and $d\in D^{\mathrm{o}}\cap U(c)\cap X$. Define $D_{\varnothing}=\overline{U(a)}\setminus U(d)$ and put $x_{0}=b$ and $x_{1}=c$. It is easy to see that (1) and (2) hold for $\sigma=\varnothing$. As for (3), $\overline{U(x_{1})}\setminus K\subset U(x_{1})\cup\alpha(x_{1})\subset U(x_{0})$.
Inductive step: Now suppose that $\sigma\in 2^{n}$ and $D_{\sigma}$ and $x_{\sigma^{\frown}0}$ and $x_{\sigma^{\frown}1}$ satisfy (1) through (3). Define
$D_{\sigma^{\frown}0}$ $=D_{\sigma}\cap\overline{V(x_{\sigma^{\frown}0})}$
$D_{\sigma^{\frown}1}$ $=D_{\sigma}\cap\overline{U(x_{\sigma^{\frown}1})}.$
By (2), Observation 6 and Lemma 5, there exists a two-sided limit point $x_{\sigma^{\frown}0,0}\in(D\cap D_{\sigma})^{\mathrm{o}}\cap V(x_{\sigma^{\frown}0})\cap X$. And there exists a two-sided limit point $x_{\sigma^{\frown}0,1}\in(D\cap D_{\sigma})^{\mathrm{o}}\cap V(x_{\sigma^{\frown}0})\cap U(x_{\sigma^{\frown}0,0})\cap X.$ Points $x_{\sigma^{\frown}1,0}$ and $x_{\sigma^{\frown}1,1}$ in $D_{\sigma^{\frown}1}$ can be found in the same way, with $U(x_{\sigma^{\frown}1})$ in the place of $V(x_{\sigma^{\frown}0})$.
Conditions (1) through (3) now hold for all $\sigma\in 2^{n+1}$. (1) is verified by combining the condition on $D_{\sigma}$ with the fact that $\overline{V(x_{\sigma^{\frown}0})}$ and $\overline{U(x_{\sigma^{\frown}1})}$ are each unions of arcs of the form $\alpha(x)$ (see Remark 2). (2) is obvious, and (3) holds for the reason given in the base case.
For all $\sigma\in 2^{<\omega}$ we now have that:
  • $\blacksquare$
    $K\cup(D_{\sigma}\cap X)$ is a connected set that meets both $K$ and $D$ by (1) and (2),
  • $\blacksquare$
    $D_{\sigma^{\frown}0},D_{\sigma^{\frown}1}\subset D_{\sigma}$ by definition, and
  • $\blacksquare$
    $D_{\sigma^{\frown}0}\cap D_{\sigma^{\frown}1}\setminus K=\varnothing$ by (3).
For every $\tau\in 2^{\omega}$ define
$K_{\tau}=\bigcap_{n=0}^{\infty}\overline{K\cup(D_{\tau\restriction n}\cap X)}.$
Then $K_{\tau}$ is the intersection of a decreasing sequence of continua stretching from $K$ to $D$, and is thus itself a continuum that meets $K$ and $D$ [12, p.6].
Figure 3. Construction of $D_{\langle 0\rangle}$ and $D_{\langle 1\rangle}$.
Figure 4. Two possible arrangements of $D_{\sigma}$ ($\sigma\in 2^{2}$) within $D_{\langle 0\rangle}$ and $D_{\langle 1\rangle}$.
Let $W$ be an open set containing $K$ and missing $D$. For each $\tau\in 2^{\omega}$ let $M_{\tau}$ be the component of some point of $K_{\tau}\cap D$ in $\overline{X}\setminus W$. Note that different $\tau^{\prime}s$ produce disjoint $M_{\tau}$’s. And by the boundary bumping principle each $M_{\tau}$ is a non-degenerate subcontinuum of $\overline{X}$ [12, Theorem 5.4]. We conclude that $\{M_{\tau}\mathrel{\mathop{\ordinarycolon}}\tau\in 2^{\omega}\}$ is an uncountable collection of pairwise-disjoint, non-degenerate subcontinua of $\overline{X}$. Therefore $\overline{X}$ is non-Suslinian.
This completes the proof of Theorem 1. $\blacksquare$

### v2

@@
 Abstract.
-We show that the closure of any self-entwined ray in the plane must contain an uncountable collection of pairwise-disjoint subcontinua. This is false in dimension three.
+We show that the closure of any self-entwined ray in the plane must contain a Cantor set of mutually disjoint continua. This is false in dimension three.
 Key words and phrases:
@@
 1. Introduction
-In this paper we investigate properties of planar continua which contain self-entwined rays.
+In this paper we study a planar form of entanglement or recurrence which gives rise to a complicated structure within the surrounding space.
+Specifically we examine properties of planar continua which contain self-entwined rays (i.e.
+rays that limit onto themselves).
+We show that these continua have uncountable structures.
 1.1. Terminology
+A ray is a continuous one-to-one image of $0,\infty)$.
+A ray is self-entwined if no open set intersects it in an arc.
+This means that each point is revisited to within $\varepsilon$ by points further along the ray (see Section 2.2).
 A continuum is a compact connected metric space.
-A ray is a continuous one-to-one image of the interval $0,\infty)$.
-A ray is self-entwined, or is a ray limiting onto itself, if no open set intersects it in an arc (see Section 2.2).
+A continuum is indecomposable if it cannot be written as a union of two proper subcontinua.
 A continuum is non-Suslinian if it contains an uncountable collection of pairwise-disjoint, non-degenerate subcontinua.
-A connected space $X$ is decomposable if there exist proper closed connected subsets $A$ and $B$ such that $X=A\cup B$; otherwise $X$ is indecomposable.
-Indecomposable continua are known to be non-Suslinian.
+Indecomposable continua are known to be non-Suslinian (cf. [8, §1.1]).
 1.2. Motivation and main result
 Self-entwined rays are commonly found in continua which arise from folding algorithms and chaotic systems.
-Such examples exist within the arc components of the Smale horseshoe, and Ikeda and Plykin attractors (see Figure 1).
-These continua are clearly non-Suslinian as they are indecomposable and even have neighborhoods homeomorphic to the Cantor fence $2^{\omega}\times[0,1]$.
+For example, they exist within the arc components of Plykin and Ikeda attractors (see Figures 1 and 2).
+These planar attractors are non-Suslinian because they are indecomposable; they even contain neighborhoods that are homeomorphic to the Cantor fence $2^{\omega}\times[0,1]$.
 The goal of this paper is to establish the general result:
 Theorem 1.
-If $X$ is a self-entwined ray in the plane, then $\overline{X}$ is non-Suslinian.
-Although every ray is a rational space, meaning that it has a basis of open sets with countable boundaries, Theorem 1 implies that the closure of a self-entwined plane ray cannot be rational.
+If $X$ is a self-entwined ray in $\mathbb{R}^{2}$, then $\overline{X}$ is non-Suslinian.
+Although every ray is a rational space, meaning that it has a basis of open sets with countable boundaries, Theorem 1 implies that the closure of a self-entwined plane ray cannot be rational (cf.
+[8, §1.3]).
 1.3. Context and limitations
-Jones [5, 6] had shown that a plane ray cannot be locally connected unless it is locally compact.
-Curry [2] proved that the closure of a self-entwined plane ray is an indecomposable continuum provided that it is $1$-dimensional and non-separating or finitely separating.
-On the other hand, it can be shown that the Sierpiński carpet, which is locally connected and $1$-dimensional, contains a dense self-entwined ray.
-There even exists a self-entwined ray in the plane that has $1$-dimensional closure, and is itself a decomposable connected set [9].
-This example shows that it is not always possible to “open up” complementary components of the closure and derive the non-Suslinian property from Curry’s result (cf.
-[4, Section 2.1]).
-Theorem 1 fails in higher dimensions by an example of Tymchatyn [13, Example 3].
-Tymchatyn’s continuum was already known to be non-planar because it is hereditarily locally connected but not finitely-Suslinian [7].
+Jones [4, 5] proved that a plane ray cannot be locally connected unless it is locally compact.
+From [7] it follows that each locally connected plane ray is homeomorphic to either $0,\infty)$, the unit circle $S^{1}$, or $S^{1}\cup[0,1]$.
+In particular, planar self-entwined rays are not locally connected.
+Tymchatyn [12, Example 3] constructed a rational continuum which is the closure of a self-entwined ray.
+So Theorem 1 is false outside the plane.
+Tymchatyn’s continuum, which can be embedded into $\mathbb{R}^{3}$, was already known to be non-planar because it is hereditarily locally connected but not finitely-Suslinian [8].
+Moreover, the ray in this continuum is locally connected.
+Curry [2] showed that the closure of a self-entwined plane ray is an indecomposable continuum provided that it is $1$-dimensional and non-separating or finitely separating.
+On the other hand, the Sierpiński carpet, which is $1$-dimensional and locally connected with infinitely-many complementary components, is easily seen to contain a dense self-entwined ray.
+So the non-Suslinian property cannot always be derived from Curry’s result.
 Figure 1.
-An Ikeda attractor.
-From [14]: “The Ikeda map is a discrete-time dynamical system that produces a strange attractor.
-It was introduced in 1979 by the physicist Kensuke Ikeda as a model for the behavior of light within a nonlinear optical resonator.
-The map demonstrates how a simple set of rules can lead to complex, chaotic behavior through a process of repeated rotation, scaling, and translation—a ‘stretch and fold’ operation common in chaotic systems.”
+The Ikeda map f(z)=1+0.9zexp(i(0.4-61+|z|2))        produces a strange attractor $K$ which contains a fixed point $p\approx 1.114-2.285i$.
+The arc component of $p$ in $K$ is a self-entwined ray with initial point $p$.
+Iterates of I={x-2.285i:0≤x≤1.114}        approximate initial segments of the ray since $f$ is injective on II, and $I\setminus\{p\}$ is contained in the basin of attraction for $K$.
+The arcs $f^{6}[I]$ and $f^{7}[I]$ are shown above.
+Figure 2.
+A planar Plykin attractor $P$ is generated by the spherical equations of [6], followed by a stereographic projection from the point $\langle 1/\sqrt{2},0,1/\sqrt{2}\rangle$.
+As with the Ikeda attractor, approximations of a self-entwined ray in $P$ are obtained by iteration on an arc $I$ which ends at a fixed point of the attractor (in planar coordinates the fixed point is $p\approx\langle 1.26,-1.53\rangle$).
+This self-entwined ray is remarkable in that it has bounded curvature.
+The images in Figures 1 and 2 were created in Python via precise numerical calculations.
 2. Preliminaries
@@
 An arc is a topological copy of the interval $[0,1]$.
-Note that if $X$ is any ray with an associated mapping $f$, and $a,b]\subset[0,\infty)$ is a closed and bounded interval, then $f\restriction[a,b]$ is a homeomorphism and $f[a,b]$ is an arc in $X$.
-It is a consequence of Sierpiński’s theorem [12, Theorem 5.16] that if $X$ is not compact then every continuum in $X$ is of this form [1, Lemma 2.2].
+If $X$ is any ray with an associated mapping $f$, and $a,b]\subset[0,\infty)$ is a closed and bounded interval, then $f\restriction[a,b]$ is a homeomorphism and $f[a,b]$ is an arc in $X$.
+By an arc of the ray $X$, we mean an arc of the form $f[a,b]$.
+It is a consequence of Sierpiński’s theorem [11, Theorem 5.16] that if $X$ is not compact then every continuum in $X$ is of this form; see [1, Lemma 2.2].
 2.2. Definition of self-entwined
-Here is an equivalent formulation of the self-entwined property in terms of the ray function.
+Below is an equivalent formulation of the self-entwined property in terms of the ray function.
 Proposition 2.
 If $X$ is a ray and $f$ is a continuous one-to-one mapping of $0,\infty)$ onto $X$, then the following are equivalent.
+  • (a)
+    $X$ is self-entwined;
+  • (b)
+    $X$ has no compact neighborhoods;
+  • (c)
     each arc of $X$ has empty interior in $X$;
+  • (d)
     for every $x\in X$ there exists a sequence $r_{n}\in[0,\infty)$ such that $r_{n}\to\infty$ and $f(r_{n})\to x$.
-Other equivalent formulations include that the space $X$ has no compact neighborhoods, or is first category in the sense of Baire.
+Proof.
+[…]
 2.3. Connectedness im-kleinen
@@
 A $\theta$-curve is a space made of three arcs $\alpha_{1},\alpha_{2},\alpha_{3}$ which have the same endpoints $a$ and $b$ and are otherwise disjoint.
-It is a standard fact of plane topology that one of the three arcs comprising a $\theta$-curve must be separated from $\infty$ by the union of the other two.
+It is a standard fact of plane topology that one of the three arcs comprising a $\theta$-curve is separated from $\infty$ by the union of the other two.
+Another fact relevant to the proposition below is that if $F$ is a closed subset of $\mathbb{R}^{2}$ which separates two points $p$ and $q$ (that is, $p$ and $q$ belong to different components of $\mathbb{R}^{2}\setminus F$), then some component of $F$ separates $p$ and $q$.
+This is a well-known corollary of Janiszewski’s theorem.
 Proposition 4.
-Let $\theta=\alpha_{1}\cup\alpha_{2}\cup\alpha_{3}$ be a theta curve of arcs joining $a$ and $b$, such that $\alpha_{1}\cup\alpha_{3}$ separates $\alpha_{2}$ from $\infty$.
+Let $\theta=\alpha_{1}\cup\alpha_{2}\cup\alpha_{3}$ be a theta curve of arcs with endpoints $a$ and $b$, such that $\alpha_{1}\cup\alpha_{3}$ separates $\alpha_{2}$ from $\infty$.
 Let $D$ be a closed disc missing $a$ and $b$.
-If $p\in\alpha_{1}\cap D$ and $q\in\alpha_{3}\cap D$, then there exists an arc $\beta\subset\alpha_{2}\cap D$ which separates $p$ from $q$ in $D$, and is such that $\beta\cap\partial D$ consists only of the two endpoints of $\beta$.
+If $p\in\alpha_{1}\cap D^{\mathrm{o}}$ and $q\in\alpha_{3}\cap D^{\mathrm{o}}$, then there exists an arc $\beta\subset\alpha_{2}\cap D$ which separates $p$ from $q$ in $D$, and is such that $\beta\cap\partial D$ consists only of the two endpoints of $\beta$.
 Sketch of proof.
-Let $\alpha_{4}$ be an arc from $a$ to $b$ in the unbounded component of $\mathbb{R}^{2}\setminus(\alpha_{1}\cup\alpha_{3})$ that misses $D$.
-Then the simple closed curve $\alpha_{2}\cup\alpha_{4}$ separates $p$ and $q$.
+Let $\alpha_{4}$ be an arc that lies in the unbounded component of $\mathbb{R}^{2}\setminus(\alpha_{1}\cup\alpha_{3})$, which has endpoints $a$ and $b$, and is such that $\alpha_{4}\cap D=\varnothing$.
+The simple closed curve $\alpha_{2}\cup\alpha_{4}$ separates $p$ and $q$.
 It follows that $\alpha_{2}\cap D$ separates $p$ from $q$ in $D$.
-Let $C$ be the component of $p$ in $D\setminus\alpha_{2}$.
-Then $\partial C$ is a simple closed curve and $\partial C\cap\alpha_{2}$ is an arc meeting the interior of $D$.
-Choose $x\in\alpha_{2}\cap\partial C\cap D^{\mathrm{o}}$ and let $\beta$ be the closure of the component of $x$ in $\partial C\cap D^{\mathrm{o}}$.
-∎
+Thus some component $\gamma$ of $\alpha_{2}\cap D^{\mathrm{o}}$ separates $p$ from $q$ in $D^{\mathrm{o}}$.
+Let $\beta=\overline{\gamma}$ (see Figure 3).∎
+Figure 3. Illustration for Proposition 4.
 2.5. Notation
-For every $A\subset\mathbb{R}^{2}$ we let $\overline{A}$, $\partial A$, and $A^{\mathrm{o}}$ denote the closure, boundary, and interior of $A$, respectively, relative to $\mathbb{R}^{2}$.
-We denote by $2^{n}$ the set of functions from $\{0,1,\ldots,n-1\}$ to $\{0,1\}$.
-Thus each $\sigma\in 2^{n}$ can be viewed as a binary sequence of length $n$.
-That is, $\sigma=\langle\sigma(0),\ldots,\sigma(n-1)\rangle$.
-We let
+For any set $A\subset\mathbb{R}^{2}$ we let $\overline{A}$, $\partial A$, and $A^{\mathrm{o}}$ denote the closure, boundary, and interior of $A$, respectively, relative to the topology of $\mathbb{R}^{2}$.
+We denote by $2^{n}$ the set of functions from $\{0,\ldots,n-1\}$ to $\{0,1\}$.
+Thus an element $\sigma\in 2^{n}$ can be viewed as a binary sequence of length $n$.
+That is, $\sigma=\langle\sigma(0),\ldots,\sigma(n-1)\rangle$ where $\sigma(i)\in\{0,1\}$ for each $i<n$.
 $2^{<\omega}=\bigcup_{n=0}^{\infty}2^{n}$
-denote the set of all finite binary sequences, and finally $2^{\omega}$ is the set of all infinite binary sequences index by $\omega=\{0,1,2,\ldots\}$.
+denotes the set of all finite binary sequences.
+Finally, $2^{\omega}$ is the set of all infinite binary sequences index by $\omega=\{0,1,2,3,\ldots\}$.
 If $\sigma\in 2^{<\omega}$, then $\sigma^{\frown}0$ and $\sigma^{\frown}1$ represent the one-element extensions of $\sigma$ obtained by appending 0 and $1$, respectively.
-If $\tau\in 2^{\leq\omega}$ and $n<\text{dom}(\tau)$, then $\tau\restriction n$ is the sequence made of the first $n$ terms of $\tau$.
-Figure 2. Illustration for Proposition 4.
+If $\tau\in 2^{\omega}$ and $n<\omega$, then $\tau\restriction n$ is the sequence made of the first $n$ terms of $\tau$.
 3. Rays that limit to arcs on both sides
-Let $X$ be a ray in $\mathbb{R}^{2}$, and let $f$ be a continuous one-to-one mapping of $0,\infty)$ onto $X$.
+Let $X$ be a ray in $\mathbb{R}^{2}$.
+Let $f$ be a continuous one-to-one mapping of $0,\infty)$ onto $X$.
 For each integer $n\geq 1$ let $S_{n}$ be a simple closed curve containing the arc $f[0,n]$.
@@
 Let $n$ be an integer such that $\beta\subset f[0,n]$.
-Since $S_{n}\cap X$ is nowhere dense in $X$, $\beta$ is contained in the union of $\overline{U_{n}\cap X}$ and $\overline{V_{n}\cap X}$.
-If each of these sets intersects $\beta$ then by connectedness of $\beta$ there exists $x\in\overline{U_{n}\cap X}\cup\overline{V_{n}\cap X}\cap\beta$.
-Then $x$ is a two-sided limit point of $X$.
+Since $\beta$ is nowhere dense in $X$, $\beta$ is contained in the union of $\overline{U_{n}\cap X}$ and $\overline{V_{n}\cap X}$.
+If each of these sets intersects $\beta$, then by connectedness of $\beta$ there exists $x\in\overline{U_{n}\cap X}\cap\overline{V_{n}\cap X}\cap\beta.$ Then $x$ is a two-sided limit point of $X$ in $U$.
 Otherwise, one of the two sets misses $\beta$ which we will show leads to a contradiction.
+The contradiction will be that there exists an arc $\gamma$ that misses $K_{2}$, and separates $p$ and $q$ in $D_{2}$.
 Without loss of generality, assume that $\overline{V_{n}\cap X}\cap\beta=\varnothing$.
 Let $G$ be the component of $\beta$ in $D_{2}\cap\overline{V_{n}}$.
-Then $G$ is a closed topological disc, $\beta\subset\partial G$, and $G\setminus\overline{V_{n}\cap X}$ is a neighborhood of $\beta$ in the space $G$.
-Thus there exists an arc $\gamma\subset G\setminus\overline{V_{n}\cap X}$ that has the same endpoints as $\beta$ and is otherwise contained in $G^{\mathrm{o}}$.
-Choosing $\gamma$ close to $\beta$ will ensure that it separates $D_{2}$ between $p$ and $q$.
-Moreover, $\gamma\cap\overline{X}\subset\partial D_{2}$ by the fact $(G^{\mathrm{o}}\setminus\overline{V_{n}\cap X})\cap\overline{X}=\varnothing$.
-Therefore $\gamma$ misses $K_{2}$, which contradicts that $K_{2}$ is connected.
+Since $\beta\cap\partial D_{2}$ consists of only the two endpoints of $\beta$, we have that $G$ is a closed topological disc and $\beta\subset\partial G$.
+Further, $G\setminus\overline{V_{n}\cap X}$ is a neighborhood of $\beta$ in the space $G$.
+Thus there exists an arc $\gamma\subset G\setminus\overline{V_{n}\cap X}$ which has the same endpoints as $\beta$ and is otherwise contained in $G^{\mathrm{o}}$.
+Choosing $\gamma$ close to $\beta$ (closer in Hausdorff distance than $d(p,\beta)$ and $d(q,\beta)$) will ensure that it separates $D_{2}$ between $p$ and $q$.
+Moreover, $\gamma\cap\overline{X}\subset\partial D_{2}$ by the fact $(G^{\mathrm{o}}\setminus\overline{V_{n}\cap X})\cap\overline{X}\subset(V_{n}\setminus\overline{V_{n}\cap X})\cap\overline{X}=\varnothing.$ Therefore $\gamma$ misses $K_{2}$, which contradicts that $K_{2}$ is connected.
 ∎
 Remark 1.
-If $\overline{X}$ is nowhere connected im-kleinen, the set of two-sided limit points of $X$ may be empty.
-Such is the case when $X$ is the accessible composant (arc component of the endpoint) of the buckethandle continuum.
+If $\overline{X}$ is indecomposable, the set of two-sided limit points of $X$ may be empty.
+This is the case when $X$ is the accessible composant (arc component of the endpoint) of the Knaster buckethandle continuum.
 4. Proof of Theorem 1
@@
 We will show that $\overline{X}$ contains an uncountable collection of pairwise-disjoint subcontinua stretching from $K$ to $D$.
-By working in the unbounded component of $\mathbb{R}^{2}\setminus K$, we may assume that if $\alpha$ is any $KDK$-minimal arc, then $\mathbb{R}^{2}\setminus(K\cup\alpha)$ has exactly two complementary components.
+By working in the unbounded component of $\mathbb{R}^{2}\setminus K$, we may assume that $K$ is non-separating.
+Hence if $\alpha$ is any $K(X\setminus K)K$-minimal arc, then $K\cup\alpha$ has exactly two complementary components.
 We may also assume that the ray’s endpoint $f(0)$ belongs to $K$, because there exists $f(t)\in K$ and $f[t,\infty)$ is a self-entwined ray with the same closure as $X$.
@@
 Note that $\alpha(x)\subset\partial U(x)\subset\alpha(x)\cup K$ and likewise for $V(x)$.
-And $\{U(x),\alpha(x),V(x)\}$ partitions the family of all $K(X\setminus K)K$-minimal arcs of $X$, with the possible exception that $\alpha(x)$ may share an endpoint in $K$ with an arc in $U(x)$ or $V(x)$.
+Additionally, $\{U(x),\alpha(x),V(x)\}$ partitions the $K(X\setminus K)K$-minimal arcs of $X$, with the possible exception that $\alpha(x)$ may share an endpoint in $K$ with an arc in $U(x)$ or $V(x)$.
 Observation 6.
@@
 $\big\{D_{\sigma}\mathrel{\mathop{\ordinarycolon}}\sigma\in 2^{<\omega}\big\},$
-and points $x_{\sigma}\in X$ so that:
+and two-sided limit points $x_{\sigma}\in X$ so that:
   1. (1)
@@
     $\overline{U(x_{\sigma^{\frown}1})}\setminus K\subset U(x_{\sigma^{\frown}0})$.
+Each $D_{\sigma}$ will be bounded by $K$ and finitely-many arcs of $X$.
 First step: To begin, let $a\in D^{\mathrm{o}}\cap X$ be a two-sided limit point of $X$ provided by Lemma 5.
@@
 Define $D_{\varnothing}=\overline{U(a)}\setminus U(d)$ and put $x_{0}=b$ and $x_{1}=c$.
-It is easy to see that (1) and (2) hold for $\sigma=\varnothing$.
-As for (3), $\overline{U(x_{1})}\setminus K\subset U(x_{1})\cup\alpha(x_{1})\subset U(x_{0})$.
+It is easy to see that (1) and (2) hold for $\sigma=\varnothing$. As for (3), $\overline{U(x_{1})}\setminus K\subset U(x_{1})\cup\alpha(x_{1})\subset U(x_{0})$.
 Inductive step: Now suppose that $\sigma\in 2^{n}$ and $D_{\sigma}$ and $x_{\sigma^{\frown}0}$ and $x_{\sigma^{\frown}1}$ satisfy (1) through (3). Define
@@
 And there exists a two-sided limit point $x_{\sigma^{\frown}0,1}\in(D\cap D_{\sigma})^{\mathrm{o}}\cap V(x_{\sigma^{\frown}0})\cap U(x_{\sigma^{\frown}0,0})\cap X.$ Points $x_{\sigma^{\frown}1,0}$ and $x_{\sigma^{\frown}1,1}$ in $D_{\sigma^{\frown}1}$ can be found in the same way, with $U(x_{\sigma^{\frown}1})$ in the place of $V(x_{\sigma^{\frown}0})$.
-Conditions (1) through (3) now hold for all $\sigma\in 2^{n+1}$.
-(1) is verified by combining the condition on $D_{\sigma}$ with the fact that $\overline{V(x_{\sigma^{\frown}0})}$ and $\overline{U(x_{\sigma^{\frown}1})}$ are each unions of arcs of the form $\alpha(x)$ (see Remark 2).
+See Figures 4 and 5.
+(1) through (3) now hold for all $\sigma\in 2^{n+1}$.
+(1) is verified by combining the condition on $D_{\sigma\restriction n}$ with the fact that $\overline{V(x_{\sigma^{\frown}0})}\cap X$ and $\overline{U(x_{\sigma^{\frown}1})}\cap X$ are unions of arcs $\alpha(x)$ (see Remark 2).
 (2) is obvious, and (3) holds for the reason given in the base case.
-For all $\sigma\in 2^{<\omega}$ we now have that:
+Figure 4. Construction of $D_{\langle 0\rangle}$ and $D_{\langle 1\rangle}$.
+Figure 5. Two possible arrangements of $D_{\sigma}$ ($\sigma\in 2^{2}$) within $D_{\langle 0\rangle}$ and $D_{\langle 1\rangle}$.
+For all $\sigma\in 2^{<\omega}$ we have:
   • $\blacksquare$
@@
 $K_{\tau}=\bigcap_{n=0}^{\infty}\overline{K\cup(D_{\tau\restriction n}\cap X)}.$
-Then $K_{\tau}$ is the intersection of a decreasing sequence of continua stretching from $K$ to $D$, and is thus itself a continuum that meets $K$ and $D$ [12, p.6].
-Figure 3. Construction of $D_{\langle 0\rangle}$ and $D_{\langle 1\rangle}$.
-Figure 4. Two possible arrangements of $D_{\sigma}$ ($\sigma\in 2^{2}$) within $D_{\langle 0\rangle}$ and $D_{\langle 1\rangle}$.
-Let $W$ be an open set containing $K$ and missing $D$.
-For each $\tau\in 2^{\omega}$ let $M_{\tau}$ be the component of some point of $K_{\tau}\cap D$ in $\overline{X}\setminus W$.
+Then $K_{\tau}$ is the intersection of a decreasing sequence of continua which contain $K$ and intersect $D$. Therefore $K_{\tau}$ is a continuum that contains $K$ and intersects $D$ [11, p.6].
+Let $W$ be an open set containing $K$ such that $\overline{W}\cap D=\varnothing$.
+For each $\tau\in 2^{\omega}$ let $M_{\tau}$ be the component of some point of $K_{\tau}\cap D$ in $K_{\tau}\setminus W$.
 Note that different $\tau^{\prime}s$ produce disjoint $M_{\tau}$’s.
-And by the boundary bumping principle each $M_{\tau}$ is a non-degenerate subcontinuum of $\overline{X}$ [12, Theorem 5.4].
+And by the boundary bumping principle each $M_{\tau}$ is a non-degenerate subcontinuum of $\overline{X}$ [11, Theorem 5.4].
 We conclude that $\{M_{\tau}\mathrel{\mathop{\ordinarycolon}}\tau\in 2^{\omega}\}$ is an uncountable collection of pairwise-disjoint, non-degenerate subcontinua of $\overline{X}$.